OpenStudy (anonymous):

find the surface area of the prism L=12IN W=12IN H=12IN

OpenStudy (anonymous):

36!

OpenStudy (anonymous):

1452 IN^2

OpenStudy (callisto):

which kind of prism? triangular? rectangular? or?

OpenStudy (anonymous):

urg

OpenStudy (anonymous):

wiat I think my first answer was correct, and yeah callisto, never thought of that....

OpenStudy (anonymous):

RECTANGULAR

OpenStudy (anonymous):

uhu, my first one was right, you just add all the sides

OpenStudy (callisto):

 you have 6 surfaces with the same area as the sides are all the same therefore surface area = (12*12)*6 = 144*6 = 864 in^2